OpenStudy (anonymous):

1) Find the value of x [-pi,pi] \[6\sin^2(2x)+5\cos[2x]=2\] 2) Find the value of x [0,2pi] 1+cosx=sinx

OpenStudy (loser66):

Where are you stuck?

OpenStudy (anonymous):

For both, I'm not sure how to start them. Any hints? @Loser66

OpenStudy (loser66):

replace sin^2 =1-cos^2 , and solve quadratic with variable is cos.

OpenStudy (loser66):

for 2) square both sides, simplify a little bit and solve for cos.

OpenStudy (anonymous):

For the second question, I got x=pi/2 and 3pi/2?? I didn't quite get the first question... @Loser66

zepdrix (zepdrix):

https://www.desmos.com/calculator/yn2gyvrbfe They clearly intersect at pi/2 and pi. When I worked out the problem I got solutions of pi/2, pi and 3pi/2. 3pi/2 is not a solution and I'm trying to understand why. I think because we did some squaring business, it may have messed things up. We have to check for extraneous solutions perhaps? Plugging 3pi/2 back into the function doesn't hold up.

zepdrix (zepdrix):

Still need some assistance on the first one? :o

OpenStudy (anonymous):

Yes, please. I need some assistance for the first question. @zepdrix

zepdrix (zepdrix):

\[\Large\rm 6\color{orangered}{\sin^2(2x)}+5\cos[2x]=2\]Applying our Square Identity to the Orange gives us,\[\Large\rm 6\color{orangered}{(1-\cos^2(2x))}+5\cos[2x]=2\]Replace cos(2x) with some variable so it's easier to work with. Say \(\Large\rm u=\cos(2x)\) So our problem is actually:\[\Large\rm 6(1-u^2)+5u=2\]You have a quadratic involving u. Solve for u by some method, factoring, quadratic formula, whatever you gotta do. You'll get some solutions, then undo the u thing. Set cos(2x) equal to those solutions, and figure out your angle from there.